OpenStudy (samigupta8):

if Sn=1+3+6+10+....+n(n+1)/2 then Sn is

OpenStudy (anonymous):

i can't think of a snap way to do it are you familiar with sigma notation? i.e. would it be okay to write this as \[\sum_{k=1}^n\frac{k(k+1)}{2}\]?

OpenStudy (samigupta8):

yaa ryt

OpenStudy (anonymous):

ok then although there might be some easier gimmick i do not see, you can write \[\sum_{k=1}^n\frac{k(k+1)}{2}=\frac{1}{2}\left (\sum_{k=1}^nk^2+\sum_{k=1}^nk\right)\]and use summation formulas to solve the ones inside the parentheses

OpenStudy (anonymous):

then a bunch of ugly algebra to "simplify" the answer

OpenStudy (anonymous):

i can walk you through it if you like but mostly it is algebra